OpenStudy (phebe):

14. Use the Zero Product Property to solve the equation: -3m(2m – 5) = 0

OpenStudy (triciaal):

if the product is zero then one factor should be zero -3m(2m-5) =0 -3m =0, 2m-5 =0 m= 0. 2m = 5 to give m = 5/2 m = 0, m= 5/2